# Interlayer exciton condensates between second Landau level orbitals in double bilayer graphene

### v1

@@ Acknowledgements.
Figure S11: Zoomed-in scans on the N = 0 interlayer states near $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,0.5)$ at $V_{\text{int}}=0$ and at 25 T. a, $R_{\text{xy}}^{\text{drag}}$. b, $R_{\text{xy}}^{\text{drive}}$. c, $R_{\text{xx}}^{\text{drive}}$. d, $R_{\text{xx}}^{\text{drag}}$. e, Dashed lines overlaid on top of the $R_{\text{xx}}^{\text{drag}}$ data, marking the strong intralayer 1/3, 2/3, 2/5, 3/5 fraction quantum Hall states in both top and bottom BLG.
@@
Here in Figs. S13a and c, we plot the same $R_{\text{xx}}^{\text{drive}}$ data as Fig. 4a and c in the main text, but as a function of $V_{\text{TG}}$ and $D_{\text{bot}}$, so we can better identify the flavors transition features in the lower BLG. They can be easily identified as the features that are generally perpendicular to $D_{\text{bot}}$ axis. Again, based on the phase diagram we obtained in Fig. 1e, here we similarly illustrate the bottom BLG LL in Figs. S13b and d, with each flavor colored following the same color scheme as in the main text, and dashed lines tracking the flavor transitions. Knowledge of the flavor distribution of the BLG confirms that the $N=1$ EC states only appear when the N = 1 wavefunction is mostly populated in the top layer of the lower BLG.

### v2

@@ Acknowledgements.
Figure S15: Zoomed-in scans on the N = 0 interlayer states near $(\nu_{\text{bot}},\nu_{\text{top}})=(0.5,0.5)$ at $V_{\text{int}}=0$ and at 25 T. a, $R_{\text{xy}}^{\text{drag}}$. b, $R_{\text{xy}}^{\text{drive}}$. c, $R_{\text{xx}}^{\text{drive}}$. d, $R_{\text{xx}}^{\text{drag}}$. e, Dashed lines overlaid on top of the $R_{\text{xx}}^{\text{drag}}$ data, marking the strong intralayer 1/3, 2/3, 2/5, 3/5 fractional quantum Hall states in both top and bottom BLG.
@@
Here in Figs. S17a and c, we plot the same $R_{\text{xx}}^{\text{drive}}$ data as Fig. 4a and c in the main text, but as a function of $V_{\text{TG}}$ and $D_{\text{bot}}$, so we can better identify the flavor transition features in the lower BLG. They can be easily identified as the features that are generally perpendicular to $D_{\text{bot}}$ axis. Again, based on the phase diagram we obtained in Fig. 1e, here we similarly illustrate the bottom BLG LL in Figs. S17b and d, with each flavor colored following the same color scheme as in the main text, and dashed lines tracking the flavor transitions. Knowledge of the flavor distribution of the BLG confirms that the $N=1$ EC states only appear when the N = 1 wavefunction is mostly populated in the top layer of the lower BLG.
@@
VII S7. Disappearance of some N = 0 interlayer states at a finite interlayer bias
Compared to Fig. 2a,b at $V_{\text{int}}=0$, the absence of N = 0 interlayer states at $(\nu_{\text{bot}},\nu_{\text{top}})\sim$ (0.5, -1.5) and (2.5, -1.5) in Fig. 3a,b at $V_{\text{int}}=0.07$ V is due to the mismatch of the orbital configuration, with one layer going into N = 1 orbital while the other remaining in N = 0. For example, at $V_{\text{int}}=0.07$ V, when top = -1.5, $D_{\text{top}}/\epsilon_{0}\sim$ 6 mV/nm. Referring to the phase diagram in Fig. 1e, this corresponds to the top BLG being in N = 1 orbital. As for the other two locations where $\nu_{\text{top}}$ = 2.5, $D_{\text{top}}/\epsilon_{0}\sim$ 120 mV/nm, which indeed corresponds to top BLG being in N = 0 orbital. However, at $(\nu_{\text{bot}},\nu_{\text{top}})$= (2.5, 2.5), the bottom layer enters the N = 1 orbital because near $\nu_{\text{bot}}$ =2.5, $D_{\text{top}}/\epsilon_{0}\sim$ -17 mV/nm, which is close to the transition between N = 0 and N = 1. In fact, this transition is clearly visible in $R_{\text{xx}}^{\text{drive}}$ as shown in Fig. S19, in the range of $2<\nu_{\text{bot}}<3$, whereas top changes, the bottom BLG orbital number oscillates between 1 and 0. This behavior appears as the alternating pattern of more resistive (white) and less resistive (dimmer) regions. For $(\nu_{\text{bot}},\nu_{\text{top}})$= (0.5, 2.5), both layers are in the N = 0 configuration, and an interlayer state is therefore expected. Indeed the characteristic $R_{\text{xx}}^{\text{drive}}$=0 diagonal feature is present (marked by the dashed square in Fig. S19). However, the corresponding drag signal is not observed at this region. While we cannot be fully certain, the proximity to the bottom BLG charge neutrality, where contact quality deteriorates, may play a role. Further systematic study with higher quality samples with improved contact quality may shed light on this observation.
Figure S19: $R_{\text{xx}}^{\text{drive}}$ same as in main text Fig. 3a. Red arrows point to where N = 1 orbital appears in the mostly N = 0 orbital in $2<\nu_{\text{bot}}<3$.